if 2 is added to half a number , the result is one more than the original number .What is the original number ?
Answer by algebrahouse.com(1659)

x = original number

x/2 + 2 = x + 1 {if 2 is added to half number, result is 1 more than number}
x + 4 = 2x + 2 {multiplied entire equation by LCD, 2, to clear fractions}
4 = x + 2 {subtracted x from both sides}
x = 2 {subtracted 2 from both sides}
original number is 2

For more help from me, visit: www.algebrahouse.com